40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 19 21:10:27 2010, yuta, Update, CDS, bad filter coefficients for every suspensions! 
    Reply  Wed Oct 20 01:22:18 2010, yuta, Update, CDS, fixed filters for MCs, but no damping Screenshot.png
Message ID: 3742     Entry time: Tue Oct 19 21:10:27 2010     Reply to this: 3744
Author: yuta 
Type: Update 
Category: CDS 
Subject: bad filter coefficients for every suspensions! 

(Koji, Yuta)

Summary:
 The damping servos of the MC suspensions has not been working since yesterday. They had worked on Friday.
 We found that some of the filter coefficients somehow changed from the correct value during dealing with the sampling frequency shift (from 2048Hz to 16384Hz). (see elog #3659)

What we did:
We thought that the problem occurred from the CDS update on Monday. So, we restored them using svn.
 1. Went to /opt/rtcds/caltech/c1/core/advLigoRTS and ran svn update using the revision number 2125.
   svn update -r 2125
 2. Rebuilt all the front ends.
   ssh -X c1sus
   cd /opt/rtcds/caltech/c1/core/advLigoRTS
   make clean-SYSNAME
   make SYSNAME
   make install-SYSNAME 
   (where SYSNAME=c1x02,c1sus,c1mcs,c1rms)
 3. Restarted the front end models.
   ssh -X c1sus
   cd /opt/rtcds/caltech/c1/scripts
   sudo startc1SYS 
   (where SYS=sus,mcs,rms)
 4. Restarted mx_streams.
   sudo /etc/restart_streams
See this wiki page for the restart procedures.

At this point we judged that the realtime system and "dataviewer" worked fine (by poking some suspensions / by measuring PSDs/TFs of the signals).

However, just restoring the RT system didn't fix the damping servos.

After that, we measured the transfer functions of the servo filters using DTT(diaggui on rosalba).
 5. The filter at MC1_SUSPOS, "3:0.0" should have a cut-off frequency at 3Hz, but it was around 0.3Hz.
 6. We used foton in order to look at the filter bank files (in /cvs/cds/rtcds/caltech/c1/chans). Then we found that the filter design was somehow changed to
   zpk([0],[0.375],2.66666,"n")
  instead of the correct one
   zpk([0],[3],0.333333,"n")

Why the filter designs changed?:
 We suspect that the filter designs were changed because of the replacement of sampling frequency in filter bank files(see elog #3659). We only replaced the lines like
  # SAMPLING SUS_MC3_SUSPOS 2048
 to
  # SAMPLING SUS_MC3_SUSPOS 16384
 and didn't changed the coefficients.
 This may caused a problem when opening the files with foton, and foton changed 3Hz to 0.375Hz (2048/16384) and other coefficients.

Note:
 The damping servo for SIDE has been working for every MCs. POS, PIT, YAW are the ones not working.

Next work:
 - Fix filters for every suspensions.
   There are backup files in /cvs/cds/rtcds/caltech/c1/chans/filter_archive directory, so this should help.
   But I don't think just fixing filters would fix the damping servo because the filter design of MC suspensions were changed this morning and the damping had worked until Friday.

ELOG V3.1.3-